OpenStudy (anonymous):

Given a = -1, h = 4, and k = -2, what is the equation of the graph if the parent function is y=√x

OpenStudy (anonymous):

Use the equation \[y=a \sqrt{x-h}+k\] Then simply plug in the numbers that they gave you.

OpenStudy (science0229):

Most textbooks show that for any function y=f(x), the transformed function is y=a f(x-h) + k So if \[f(x)=\sqrt{x}\], the transformed form would be \[-\sqrt{x-4}-2\]

OpenStudy (anonymous):

@Nekojellehfish Does this make sense?

OpenStudy (anonymous):

@RPguy Yeah thanks a lot :D

OpenStudy (anonymous):

@Nekojellehfish No problem :)